### v1

# From ABC to Effective Roth and Ridout Constants for Cubic Roots

Abstract
Enrico Bombieri showed conditionally (1994) that the ABC conjecture implies Roth’s theorem, and Van Frankenhuysen (1999) later provided a complete proof. Building on Bombieri’s and Van der Poorten’s explicit formula for continued-fraction coefficients of algebraic numbers—specialized to cubic roots—we derive an effective bound for a Roth-type constant assuming an effective form of ABC. Roth’s original argument establishes existence but does not yield an explicit value; our approach makes the dependence on the ABC parameters explicit and also gives an explicit bound in the corresponding special case of Ridout’s theorem.
We then introduce the notion of approximation gain as a refinement of the “quality” of an abc-triple. For c in a large computational range (verified up to $c<2^{63}$), the approximation gain remains below a strikingly small threshold, motivating the conjecture that the approximation gain is always smaller than $\frac{3}{2}$. This suggests a potential strategy for attacking ABC by bounding approximation gain and power gain separately.  [^1] The authors are grateful to Benne de Weger, Joshua Lampert, Timm Lampert, Enrico Bombieri, Wadim Zudilin, Michel Waldschmidt, Preda Mihailescu, Paul Vojta, Stephane Fischler, Robin Zhang, Abderrahmane Nitaj, Sam Chow, Noah Lebowitz-Lockard, Ingo Althöfer, ChatGPT and Ulrich Tamm for helpful comments.
1 Introduction
Roth’s theorem is famously deep, and its original proof is both long and ineffective (see Roth (1955)). In particular, it guarantees the existence of a constant in the Diophantine approximation bound, but it does not provide an explicit value or a practical way to compute one.
Bombieri proved conditionally in 1994 that Roth’s theorem follows from the ABC conjecture, and Van Frankenhuysen later gave a full proof in 1999 (Van Frankenhuysen (1999)). In this paper we revisit that implication with an additional goal: making the constants explicit in a natural family of cases. We focus on cubic roots and use an explicit continued-fraction formula of Bombieri and Van der Poorten (Bombieri &\& Van der Poorten (1975)) for the coefficients of regular continued fractions of algebraic numbers. This yields a concrete relationship between the $\varepsilon$-parameters in Roth-type inequalities and the corresponding $\varepsilon$ in ABC, and it leads to an explicit upper bound for the inverse of the Roth constant in our setting.
In our earlier paper (Sibbertsen (2022)), we showed that a weakened form of ABC follows from Roth’s theorem in certain special cases. The present work uses essentially the same vocabulary and techniques, but in the opposite direction: assuming (an effective form of) ABC, we extract explicit Diophantine approximation bounds.
We also introduce the approximation gain, which is always lower or equal to the quality of an ABC hit.
2 Approximation gains for ABC
2.1 Roth’s theorem
In the following, the coefficients of the regular continued fraction are called $b_{n}$ and the approximants $\frac{p_{n}}{q_{n}}$.
Theorem 1
Roth’s Theorem: Let $a$ be an algebraic number. Then for every $\varepsilon>0$, there exists a constant C dependent on the algebraic number $a$ and $\varepsilon$ such that for all positive integers $p$ and $q$:
$|a-\frac{p}{q}|>\frac{C}{q^{2+\varepsilon}}.$ (1)
Proof 1
See Roth (1955).
Van Frankenhuysen (1999) discusses Roth’s Theorem and its implication by the ABC conjecture in detail and investigates effective versions of Mordell’s conjecture. Furthermore, Granville and Tucker (2002, p.1229) show that a generalized version of Roth’s theorem implies the ABC conjecture effectively.
With the help of the inequality for regular continued fractions from this follows $b_{n+1}\leq\frac{q_{n}^{\varepsilon}}{C}$
2.2 Formula by Bombieri and Van der Poorten
Our main technical input is the explicit formula of Bombieri and Van der Poorten for the continued-fraction coefficients of algebraic numbers by (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)), in whichf(x){f(x)} is the minimal polynomial of the algebraic number.
The formula links two quantities—call them $d_{n}$ and $b_{n}$ — and is therefore well suited to bridge ABC-type information (which naturally arises from integer identities involving $d_{n}$ with Diophantine approximation data encoded by $b_{n}$).
$\frac{f^{\prime}(x)}{f(x)}$ leads to $d_{n}$ in the denominator. This holds because $\frac{f^{\prime}(x)}{f(x)}$ results in $\frac{f^{\prime}(\frac{p_{n}}{q_{n}})}{f(\frac{p_{n}}{q_{n}})}=\frac{q_{n}^{s}sp_{n}^{s-1}}{d_{n}q_{n}^{s-1}}$ for the convergents with $f(x)=x^{s}-k$ being the minimal polynomial of the $s$th root of k.
For the cubic case the error term of the formula (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)) can be ignored due to the effective bound by Liouville’s theorem.
2.3 ABC Conjecture
To formulate the ABC Conjecture we need the following definition:
Definition 1
For a positive integer $a$, $rad(a)$ is the product of the distinct prime factors of $a$.
Definition 2 (ABC Conjecture)
For every positive real number $\varepsilon$, there exists a constant $K_{\varepsilon}$ such that for all triples $(a,b,c)$ of coprime positive integers, with $a+b=c$:
$c<K_{\varepsilon}\cdot rad(abc)^{1+\varepsilon}.$
Note that this version of the ABC Conjecture is not effective. This is so because only the existence of a constant $K_{\varepsilon}$ is claimed without specifying how to calculate $K_{\varepsilon}$ for given $\epsilon$.
2.4 Resulting ABC equations
We look at 3rd roots of an integer $k$ which is not a perfect cube. The coefficients of the regular continued fraction are called $b_{n}$ and the approximants $\frac{p_{n}}{q_{n}}$.
We call the equations with positive numbers
$p_{n}^{3}=kq_{n}^{3}+d_{n}$
resp
$kq_{n}^{3}=p_{n}^{3}+d_{n}$
resulting equations.
To use ABC we have to get rid off the cofactors of the resulting equations. For this purpose we observe that gcd($kq_{n}^{3},d_{n}$) is the natural greatest common divisor to divide out.
This will be called gcd in the following. It is larger or equal to 1 and smaller or equal to k as gcd($q_{n},p_{n}$) = 1.
So with
$\frac{p_{n}^{3}}{gcd}=\frac{kq_{n}^{3}}{gcd}+\frac{d_{n}}{gcd}$
we get an ABC equation.
2.5 Construction of the effective bound for the inverse of Roth’s constant
Theorem 2
From an effective version of the ABC conjecture Roth’s Theorem effective for resulting equations of 3rd roots follows: Let $a$ be the 3rd root of k. Then for every $\varepsilon>0$, there exists a constant C dependent on the algebraic number $a$ and $\varepsilon$ such that for all positive integers $p$ and $q$:
$|a-\frac{p}{q}|>\frac{C}{q^{2+\varepsilon}}.$ (2)
And the inverse of Roth’s constant C is smaller or equal to $K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{1}}{q_{1}})^{\frac{3\varepsilon}{1+\varepsilon}}$
Proof: Let $\varepsilon$ be the ABC $\varepsilon$ and $K_{\varepsilon}$ the ABC constant.
The ABC conjecture implies
$\frac{p_{n}^{3}}{gcd}\leq K_{\varepsilon}Rad(p_{n}q_{n}k(d_{n}gcd^{-1}))^{1+\varepsilon}$
This impiles
$p_{n}^{3}\leq K_{\varepsilon}(p_{n}q_{n}d_{n}k)^{1+\varepsilon}\frac{gcd}{gcd^{1+1\varepsilon}}$
In the cubic case, Bombieri–Van der Poorten’s explicit continued-fraction formula (Bombieri and Van der Poorten (1975), Theorem 3, formula (13)) applies without the final error term, thanks to effective Liouville bounds; compare also their worked example for $\sqrt[3]{2}$ (loc. cit., p. 152). In our case their algorithm is also always reduced after the first step.
For $n>=4$ an error term can arise and the method can not always be started with the first step. This is the reason, why we restrict ourselves to 3rd roots here.
From (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)), we have the inequality
$d_{n}\leq\frac{3p_{n}^{2}}{q_{n}b_{n+1}}$
Using that inequality we get
$b_{n+1}\leq K_{\varepsilon}^{1/(1+\varepsilon)}3kp_{n}^{\frac{3\varepsilon}{1+\varepsilon}}({\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
Roth’s theorem implies the existence of a constant C depended only on $\varepsilon_{Roth}$ and 3rd root of k with
$b_{n+1}\leq\frac{q_{n}^{\varepsilon_{Roth}}}{C}$
So as relationship between the epsilons we get
$\varepsilon_{Roth}=3(\frac{\varepsilon_{ABC}}{1+\varepsilon_{ABC}})$
and we get that the inverse of Roth’s constant C can be bounded by
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{n}}{q_{n}})^{\frac{3\varepsilon}{1+\varepsilon}}({\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
Now the gcd term is smaller or equal to 1 and $\frac{p_{n}}{q_{n}}$ is bounded in this case from above by $\frac{p_{1}}{q_{1}}$ so the desired result follows. All following convergents are smaller.
2.6 Range of the Epsilons
For $\varepsilon_{ABC}$ = $\varepsilon_{Roth}$ = 0 it results that $b_{n}$ is bounded. This is an open question, but it is known that ABC is not valid for $\varepsilon=0$, but the known counterexamples are not resulting equations of 3rd roots.
For $\varepsilon_{ABC}$ = $\varepsilon$ = $\frac{1}{2}$ we get $b_{n+1}\leq q_{n}\frac{p_{n}}{q_{n}}factor$, which bascially is Liouville’s theorem. So the interesting part is $\varepsilon_{ABC}$ between 0 and $\frac{1}{2}$.
2.7 Example $\sqrt[3]{2}$
The regular continued fraction starts with $[1;3,1,5,..]$
The first resulting equations are
2 = 1 + 1, which is an ABC equation
64 = 54 + 10, where $gcd=2$ so we get the ABC equation $32=27+5$
128 = 125 + 3, which is an ABC equation
Our formula for a bound of the inverse of Roth’s constant
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{n}}{q_{n}})^{\frac{3\varepsilon}{1+\varepsilon}}{(\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
leads in this case to
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}6(\frac{4}{3})^{\frac{3\varepsilon}{1+\varepsilon}}$
as $\frac{4}{3}$ is the largest approximant and the factor with $gcd=1$ is larger than the one with $gcd=2$, so that it can be taken and is equal to 1.
There is a lot of data on ABC, which could be used to get a bound for the inverse of Roth’s constant within the range already analysed by computer calculation. We will investigate $\sqrt[3]{2}$ now for 2 different $\varepsilon$.
2.8 $\varepsilon=0.5$
Due to Korobov’s famous result Korobov (1990)
$\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{1}{q^{2.5}}$
for all natural numbers $p$ and $q$ with the exception of 1 and 4 for $q$
If we ignore 1 then Roth’s constant is calculated by
$\mid\sqrt[3]{2}-\frac{5}{4}\mid 4^{2.5}$
which is approximately 0,32 and so the the inverse is given by
$\frac{1}{C}=3,125$ approximately.
Now we use our way from the ABC side. A $\varepsilon_{Roth}=0.5$ leads to an $\varepsilon_{ABC}=\frac{1}{5}$
Again ignoring the first case 2 = 1 + 1 we take from the ABC data the $K_{\varepsilon}$ that results from the resulting equation 128 = 125 + 3.
So calculation gives $K_{\frac{1}{5}}$ is approximately 2,16
So our formula for a bound of the inverse of Roth’s constant results in
$\frac{1}{C}\leq 2,16^{\frac{5}{6}}6(\frac{4}{3})^{\frac{1}{2}}$
which is approximately 13,16
So in this case the road via ABC leads to a loss in precision, which is mainly due to the use of $\frac{4}{3}$ as upper bound.
2.9 $\varepsilon=0.4$
Here the result is not known exactly. In the proof of corrollary 2.2 Voutier (2007) gives one result using the hypergeometric method for $\varepsilon=0.4325$ using his theorem 2.1 (p.285):
$\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{10^{-99}}{q^{2.4321}}$
So the inverse of Roth’s constant is bounded by $10^{99}$.
Now we use our way from the ABC side and use a $\varepsilon_{Roth}=0.4$. This leads to an $\varepsilon_{ABC}=\frac{2}{13}$
Again ignoring the first case 2 = 1 + 1 from the ABC data we take the ABC $K_{\varepsilon}$ that results from the resulting equation 128 = 125 + 3 in the following.
So calculation gives $K_{\frac{2}{13}}$ is approximately 2,527
So our formula for a bound of the inverse of Roth’s constant results in
$\frac{1}{C}\leq 2,527^{\frac{13}{15}}6(\frac{4}{3})^{\frac{6}{15}}$
which is approximately 15,03
2.10 Table for bounds of the inverse of Roth’s constant C
We conclude with a table. Again we ignore the resulting equation 2 = 1 + 1 and assume that the resulting $K_{\varepsilon}$ comes from the resulting equation 128 = 125 + 3. This is not always the case as will become clear from the data. Only with $\varepsilon$ large enough in the known data range it comes from that equation.
$\varepsilon_{Roth}$ $\varepsilon_{ABC}$ $K_{\varepsilon}$ ABC bound for $\frac{1}{C}$ Known bound for $\frac{1}{C}$
0 0 4,267 60,686 Probably unbounded
0,4 $\frac{2}{13}$ 2,5 15,03 Not known
0,5 $\frac{1}{5}$ 2,161 13,16 3,125 (Korobov’s result is exact)
1 $\frac{1}{2}$ 0,78 6,78 6 (Liouville can be lowered to 1,575)
Table 1: Approximate values for $\sqrt[3]{2}$ as long as this resulting $K_{\varepsilon}$ can be taken
In general the bound resulting from this ABC method is very smooth, but not as good as the exact results of course. Of course if ABC is proved for an explicit $\varepsilon$ then that $K_{\varepsilon}$ can be taken and no numerical data would be needed. At the moment one open conjecture is that ABC holds with $\varepsilon$ = 1 and $K_{\varepsilon}$ = 1. But to get results for Roth’s constant C a much smaller $\varepsilon$ would be needed.
2.11 An explicit bound for Ridout’s theorem
Using the ABC method it is also possible to give an explicit bound for the number of solutions of Ridout’s theorem for the special case of the approximants of the regular continued fraction:
Theorem 3
(D. Ridout, 1957):
Let S be a finite set of prime numbers. For any real algebraic number a, for any $\varepsilon>0$, the set of $\frac{p}{q}$ with p an integer and q a S–integer and
$\mid a-\frac{p}{q}\mid<{q^{-1-\varepsilon}}$
is finite.
See Waldschmidt (2008) page 75.
2.12 Square Roots
As we are dealing with the prime factorisation, here the squares roots are interesting as well and a bound follows from ABC as follows:
For $\sqrt{k}$ the resulting equations are
$p_{n}^{2}=kq_{n}^{2}+1$
resp
$kq_{n}^{2}=p_{n}^{2}+1$
We only look at the first case here. The second is basically analogous. Here the gcd is always 1 and so we have ABC equations and get from ABC
${p_{n}^{2}}\leq K_{\varepsilon}Rad(p_{n}q_{n}k)^{1+\varepsilon}$
From this follows
${p_{n}^{2}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{1+\varepsilon}p_{n}^{1+\varepsilon}$
and so with $Rad(q_{n})$ given as $q_{n}$ is an S integer we get the bound
${p_{n}^{\frac{2}{1+\varepsilon}}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{1+\varepsilon}$
for the $p_{n}$ and so the number of solutions must be finite and an explicit bound is given, which follows from an explicit ABC version.
We assume now that the ABC conjecture is valid for $\varepsilon=1$} with $K_{\varepsilon}=1$. Then for $\sqrt{2}$ results
prime numbers Bound for $p_{n}$ Possible Approximants
2 16 $\frac{3}{2}$}
3 36 None as $\frac{4}{3}$} is not in the regular continued fraction
5 100 $\frac{7}{5}$}
2*3 144 $\frac{3}{2}$} and $\frac{17}{12}$}
Table 2: The first values for $\sqrt{2}$
2.13 Third Roots
Now we use our method from above based on the formula of Bombieri and Van der Poorten.
For the third root of k the resulting equations are
$p_{n}^{3}=kq_{n}^{3}+d_{n}$
resp
$kq_{n}^{3}=p_{n}^{3}+d_{n}$
We only look at the first case here. The second is basically analogous. Here the gcd is not always 1 and a priori must be considered. But in the resulting inequality the case with gcd=1 is the critical one so we can assume that from the beginning. So in this sense we have ABC equations and get from ABC
${p_{n}^{3}}\leq K_{\varepsilon}Rad(p_{n}q_{n}d_{n}k)^{1+\varepsilon}$
For 3rd roots the formula (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)) by Bombieri and Van der Poorten is valid without error term due to Liouville’s theorem as given by Bombieri and Van der Poorten and their method can be used directly. For $n>=4$ an error term can arise and the method can not be started with the first step. This is the reason, why we restrict ourselves to 3rd roots here.
From (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)), we have the inequality
$d_{n}\leq\frac{3p_{n}^{2}}{q_{n}b_{n+1}}$
Using that inequality we get
${p_{n}^{3}}\leq K_{\varepsilon}Rad(p_{n}q_{n}d_{n}k)^{1+\varepsilon}$
${p_{n}^{3}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}(p_{n}d_{n}k)^{1+\varepsilon}$
${p_{n}^{3}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}(p_{n}k)^{1+\varepsilon}(\frac{3p_{n}^{2}}{q_{n}b_{n+1}})^{1+\varepsilon}$
And so using $b_{n+1}>=1$} finally
${p_{n}^{\frac{3}{2(1+\varepsilon)}}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{2+2\varepsilon}(\frac{p_{n}}{q_{n}})^{1+\varepsilon}$
for the $p_{n}$ and so the number of solutions must be finite and an explicit bound can be given as $\frac{p_{n}}{q_{n}}$ converges to the 3rd root of k. So one choice is
$p_{n}\leq(K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{2+2\varepsilon}(\frac{p_{1}}{q_{1}})^{1+\varepsilon})^{\frac{2({1+\varepsilon})}{3}}$
2.14 Bounds for the approximation gain of ABC hits from resulting equations
Definition 3
A triple of coprime integers $a+b=c$ is an ABC hit, if $rad(abc)<c$ holds and the quality of a triple is given by the first $quality(a,b,c)=\frac{ln(c)}{ln(rad(abc))}$. As $c>1$ it is clear that abc is larger than 1.
$2+23^{5}=9^{5}\cdot 109$ is the ABC hit with the highest quality so far found. Its quality is approximately 1.62991. This hit results from the 3rd convergent to $\sqrt[5]{109}$ with regular continued fraction $2;1,1,4,77733,\ldots]$. The fraction $\frac{23}{9}$ is a good approximation because 77733 is very large and very near the absolute Liouville bound.
Definition 4
Let $d_{n}q_{n}kp_{n}$ come from a resulting equation of $\sqrt[s]{k}$. The approximation gain of (a,b,c) equals $\frac{ln({q_{n}}^{s}k)}{ln(d_{n}q_{n}kp_{n})}$ if $d_{n}>0$, and $\frac{ln({p_{n}}^{s})}{ln(-d_{n}q_{n}kp_{n})}$ if $d_{n}<0$.
The approximation gain is always smaller or equal to the quality.
Theorem 4
For resulting equation from 3rd roots we now have that the approximation gain is always smaller than $\frac{3}{2}$, if $p_{n}>q_{n}$.
Proof 2
We need two inequalities. Liouville’s bound $b_{n+1}\leq 3kq_{n}$
and from (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)), we have the inequality
$d_{n}\geq\frac{3p_{n}^{2}}{q_{n}(b_{n+1}+2)}$
So we get approximation gain = $\frac{ln({p_{n}}^{3})}{ln(-d_{n}q_{n}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{q_{n}(b_{n+1}+2)}q_{n}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{(b_{n+1}+2)}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{(3kq_{n}+2)}kp_{n})}\leq$ $\frac{3}{2}\frac{ln({p_{n}})}{ln(\frac{3kp_{n}}{(3kq_{n}+2)})}<$ $\frac{3}{2}$
When it is assumed that the method of Bombieri and Van der Poorten works as well for $\sqrt[m]{k}$ with $m\geq 3$ and $k\geq 2$ then the following would result:
The inequality of Bombieri and Van der Poorten is $d_{n}\geq\frac{mp_{n}^{m-1}}{q_{n}(b_{n+1}+2)}$
Liouville’s bound is $b_{n+1}\leq mkq_{n}$
And the resulting inequality is
approximation gain = $\frac{ln({p_{n}}^{m})}{ln(-d_{n}q_{n}kp_{n})}\leq$ $\frac{m}{2}$
So even when it would work (the error terms can be ignored and the algorithm started, which is clear for 3rd degree) the bound for this inequality would diverge and so no universal bound for all m can be found on this road as Liouville’s bound is not strong enough for this purpose.
So the ABC gains can be seperated for resulting equation in approximation gains and power gains. The approximation gain can be bounded by $\frac{3}{2}$ for 3rd roots. To prove a bound for the quality the power gains $\frac{ln({-p_{n}q_{n}d_{n}k})}{ln(rad(-d_{n}q_{n}kp_{n}))}$ must be bounded as well, which seems to be very difficult even for special cases like $\sqrt[3]{2}$ (remarkable are here the second equation 128 = 125 + 3 with a power gain of 1,4 and an approximation gain of 1,01 and the 5th equation with a power gain of 1,37 and an approximation gain of 0,99).
To bound the power gains a strong explicit effective ABC result is needed. There is a discussion in the mathematical community, if Mochizuki’s results can be used. We do not know this, but if they can be used the following follows:
Theorem 5 (Explicit ABC)
For every positive real number $\varepsilon$, there exists a constant $L_{\varepsilon}$ such that for all triples $(a,b,c)$ of coprime positive integers, with $a+b=c$:
$c<L_{\varepsilon}\cdot rad(abc)^{3+\varepsilon}.$
This can be used to prove that the power gains of the quality for resulting ABC equations of 3rd roots are bounded for a fixed k:
power gain = $\frac{ln({-p_{n}q_{n}d_{n}k})}{ln(rad(-d_{n}q_{n}kp_{n}))}$
$\leq\frac{ln({k}^{2}{p_{n}}^{3})}{ln(rad(-d_{n}q_{n}kp_{n}))}$
using the formula (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)) by Bombieri and Van der Poorten and the fact that $b_{n+1}>=1$.
= $\frac{ln({k}^{2}{p_{n}}^{3})}{ln((rad(-d_{n}q_{n}kp_{n})^{3+\varepsilon})^{\frac{1}{3+\varepsilon}})}$
$\leq\frac{ln({k}^{2}{p_{n}}^{3})}{ln(\frac{{p_{n}}^{3}}{L_{\varepsilon}}^{\frac{1}{3+\varepsilon}})}$
for n large enough. This converges to $3+\varepsilon$} and is therefore bounded for all n. The bound itself depends on k and is of course extremely large.
Recently Zhou (2025) has reduced the constant $L_{\varepsilon}$ from about $10^{30}$ to 400 under certain extra conditions.
One concept to prove the ABC conjecture might then be to prove that both gains are bounded seperately. We conjecture that the approximation gain is bounded by $\frac{3}{2}$.